# An Exact Solution Algorithm for the Bi-Level Optimization Problem of Electric Vehicles Charging Station Placement

### v1

@@ 3.1 LL Problem (User Charging Decision)
In real networks, not all EV travel demands between OD pairs can be satisfied due to limited battery range and sparse charging infrastructure. This is an issue for travel over long distances, since usually the trips within a city can be completed without needing to charge along the route. As a result, for trips with long distances, certain trips become infeasible if no reachable path exists that either stays within the battery limit or includes a charging opportunity. This constraint motivates defining the set of feasible paths based on battery limitations, which acknowledge that some OD pairs may have zero flow in the UE solution. The feasible paths available to each user depend on the installed charging infrastructure $\mathbf{y}$.
@@
The LL problem thus jointly determines equilibrium path flows, link flows, and charging station demand, conditional on the UL decision $\mathbf{y}$. We consider a network $\mathcal{G}=(\mathcal{N},\mathcal{A})$: graph $\mathcal{G}$ is the tuple $(\mathcal{N},\mathcal{A})$ comprised of the set of nodes $\mathcal{N}$ and the set of links $\mathcal{A}$. Each link can be denoted by a pair of nodes $(i,j)$. Let $\mathcal{Z}\subseteq\mathcal{N}$ be the set of zones, which represent nodes that are the origins or destinations of the travelers. $\mathcal{C}\subseteq\mathcal{N}$ is the set of candidate locations for building charging infrastructure. A path is feasible for an EV if either the usage of the battery is less than the battery limit, assuming that users will start their trip with a full battery, or there is a charging station on the route where the user recharges the battery to reach the destination before the battery is depleted. Therefore, not all demand from an origin to a destination can be met, and it depends on the $\mathbf{y}$, which specifies where charging stations are located and consequently which paths are feasible. Let $\Pi_{rs}(\mathbf{y})$ be the set of feasible paths based on the charging stations location $\mathbf{y}$ between the origin $r\in\mathcal{Z}$ and destination $s\in\mathcal{Z}$. Let $\Pi(\mathbf{y)}=\cup_{(r,s)\in\mathcal{Z}^{2}}\Pi_{rs}(\mathbf{y})$ be the set of all feasible paths. Travel time is considered the travel cost. The travel time for the link $(i,j)$ is based on a convex function of traffic flow on this link, which is $t_{ij}(x_{ij})$, where $x_{ij}$ is the link $(i,j)$ flow. The full formulation of this convex UE problem is given below.
$\min_{\mathbf{x},\mathbf{h},\mathbf{v}}\quad$ $L(\mathbf{y,x(h),v(h)})=\sum_{(i,j)\in\mathcal{A}}\int_{0}^{x_{ij}}t_{ij}(x)\,dx+\sum_{l\in\mathcal{C}}\int_{0}^{v_{l}}t\,dv$ (1a)
@@
$v_{l}=\sum_{\pi\in\Pi(\mathbf{y})}\gamma_{l}^{\pi}h_{\pi},$ $\quad\forall l\in\mathcal{C}$ (1e)
@@
The LL problem (1) constitutes a modified traffic assignment problem (MTAP) that accounts for the unique cost structure faced by EV users. In contrast to classical TAP formulations, which only consider travel time as the cost, our model includes both travel time and charging costs, which indicates EV users must also select charging stations along their routes. This extension is reflected in the objective function, which includes an additional term for the cost of charging based on the station prices and the flow of users choosing each station. Furthermore, two additional constraints distinguish this model from standard TAP: (i) a station activation constraint (1b) ensures no users can charge at the location $l$ unless a charging facility is built there. In this constraint, $y_{l}$ comes from UL and is fixed here, and (ii) a flow conservation constraint that determines the amount of charging demand at each station based on users’ path choices. These modifications are grounded in the assumption that EV travelers behave rationally under perfect information and choose paths that minimize their total travel and charging cost.
We will name this problem MTAP ($\mathbf{y}$), and let $L(\mathbf{y},\mathbf{x}(\mathbf{h}),\mathbf{v}(\mathbf{h}))$ be the objective function value. To achieve the travel time between $r$ and $s$, we need to sum up all the links’ travel time between $r$ and $s$, which are in the user’s path. The link flows, and the path flows, denoted as $h_{\pi}$, are closely related. Let $\delta^{\pi}_{ij}$ denote whether link $(i,j)$ is used by path $\pi$, so $\delta^{\pi}_{ij}=0$ if path $\pi$ of mode $m$ does not use link $(i,j)$, and $\delta^{\pi}_{ij}=1$ if it does. Based on constraint (1c), link flows are the sum of all the path flows that use that link.
@@
Let $v_{l}$ be the total time spent charging at station $l$ per hour aggregated over all EV users who stop there. In other words, $v_{l}$ represents the total time EV users collectively spend charging at station $l$ per hour. We can only have a positive value for $v_{l}$ if the decision is to build a charging station there. Constraint (1b) shows this linking relation between $v_{l}$ and $y_{l}$. So if $y_{l}$ is 0, then $v_{l}$ must be 0 too. In constraint (1e), $\gamma^{\pi}_{l}$ is the amount of time (in hours) that a vehicle traveling along path $\pi$ must spend charging at station $l$, and $h_{\pi}$ is the flow on that path. Therefore, $v_{l}$ is equal to the sum of total charging time per unit time spent at station $l$, where $l$ is one of the nodes of that path and $l$ is not the origin or destination of path $\pi$. We have assumed that the total charging cost for each user depends on their battery level prior to charging, and there is no congestion or delay at the charging stations, because our model implicitly assumes that stations are designed with adequate capacity to meet demand. In practice, this assumption can be satisfied by proper sizing of the charging infrastructure (that is, installing sufficient charging ports) based on the demand patterns identified by our model. Our problem focuses on strategic placement, not detailed operational dynamics. Queuing effects and congestion are second-order phenomena that can be addressed in a separate operational model.
The constraints (1g) and (1f) require the non-negativity for the flow of the charging station and the flow of the path. The first term in the objective function of problem (1) corresponds to the classical Beckmann transformation, which is a potential function whose minimizer satisfies the UE conditions with respect to travel time. The second term incorporates the charging cost and extends the potential function to reflect users’ sensitivity to both travel and charging components in their path selection. While the total objective does not represent a cost directly minimized by users, its minimization yields a flow pattern consistent with UE behavior under combined travel and charging costs. $\lambda_{l}^{\pi}$ is the time spent charging on $l$ coming from path $\pi$. The total time spent on charging $v_{l}$ will be multiplied by $t$, which is a fixed price that accounts for both the charging cost and the time spent charging.
@@ 4.1 High Point Relaxation (Global Lower Bound)
We will derive the HPR problem by considering only the follower’s constraints without taking into account its objective function.
@@
$y_{l}\in\{0,1\},$ $\quad\forall l\in\mathcal{C}$ (4c)
@@
$v_{l}=\sum_{\pi\in\Pi(\mathbf{y})}\delta_{l}^{\pi}h_{\pi},$ $\quad\forall l\in\mathcal{C}$ (4g)
@@ 4.2 Value Function Cuts
$y_{l}\in\{0,1\},$ $\quad\forall l\in\mathcal{C}$ (6c)
@@
$v_{l}=\sum_{\pi\in\Pi(\mathbf{y})}\delta_{l}^{\pi}h_{\pi},$ $\quad\forall l\in\mathcal{C}$ (6g)
@@
$L(\mathbf{y},\mathbf{v}(\mathbf{h}),\mathbf{x}(\mathbf{h}))\leq L(\mathbf{y},\mathbf{v}^{\prime}(\mathbf{h}^{\prime}),\mathbf{x}^{\prime}(\mathbf{h}^{\prime})),$ $\quad\forall\,(\mathbf{v}^{\prime}(\mathbf{h}^{\prime}),\mathbf{x}^{\prime}(\mathbf{h}^{\prime}))\in\mathcal{H}(\mathbf{y})$ (6j)
Constraint (6j) represents the value function cuts. $\mathcal{H}(\mathbf{y})$ represents the set of possible $(\mathbf{v}^{\prime}(\mathbf{h}^{\prime}),\mathbf{x}^{\prime}(\mathbf{h}^{\prime})$ under $\mathbf{y}$ design. In particular, for each $\mathbf{v}^{\prime}(\mathbf{h}^{\prime}),\mathbf{x}^{\prime}(\mathbf{h}^{\prime})\in\mathcal{H}(\mathbf{y})$, the derived aggregate flows $\mathbf{v}^{\prime}(\mathbf{h}^{\prime})$ and $\mathbf{x}^{\prime}(\mathbf{h}^{\prime})$ respect $v_{l}^{\prime}=0$ if $y_{l}=0$ for all $l\in\mathcal{C}$, so that no site without an established charging station is assigned any charging demand. By restricting to flow configurations consistent with the given $\mathbf{y}$, this formulation avoids infeasible cases such as $v_{l}^{\prime}>0$ when $y_{l}=0$.
@@
It is sufficient to show equality of the feasible regions, since the objective functions in both problems are identical. Suppose $(\mathbf{y},\mathbf{v},\mathbf{x})$ is feasible for problem (6). This means $(\mathbf{y},\mathbf{v},\mathbf{x})$ satisfies the value function cuts in constraint (6j). Therefore, under the same $\mathbf{y}$, the $L(\mathbf{y},\mathbf{v}(\mathbf{h}),\mathbf{x}(\mathbf{h}))$ is less than or equal to the $L(\mathbf{y},\mathbf{v}^{\prime}(\mathbf{h}^{\prime}),\mathbf{x}^{\prime}(\mathbf{h}^{\prime}))$ $\forall\,(\mathbf{v}^{\prime}(\mathbf{h}^{\prime}),\mathbf{x}^{\prime}(\mathbf{h}^{\prime}))\in\mathcal{H}(\mathbf{y})$. This directly means $(\mathbf{v},\mathbf{x})$ minimizes $L(\mathbf{y},\mathbf{v}(\mathbf{h}),\mathbf{x}(\mathbf{h}))$ given the decision $\mathbf{y}$. Thus, we can ensure the solution is feasible for the original bi-level problem (6). Conversely, consider any point $(\mathbf{y},\mathbf{v},\mathbf{x})$ feasible for the original bi-level problem (6). By the definition of LL optimality, $(\mathbf{v},\mathbf{x})$ minimizes $L(\mathbf{y},\mathbf{v}(\mathbf{h}),\mathbf{x}(\mathbf{h}))$ given the decision $\mathbf{y}$. Therefore, constraint (6j) is trivially satisfied, implying feasibility for problem (6). Hence, both problems have identical feasible sets and are thus equivalent. ∎
@@
While value function cuts are useful for iteratively refining the feasible region of the HPR, their applicability is conditioned on the feasibility of the follower solution under the current UL decision $\mathbf{y}$. In particular, a previously generated pair $(\mathbf{v}^{\mathrm{f}},\mathbf{x}^{\mathrm{f}})\in\text{MTAP}(\mathbf{y}^{\prime})$ may no longer be valid if certain charging stations used in that solution are deactivated in the new $\mathbf{y}$. To prevent invalid comparisons and over constraining the feasible region, we introduce a relaxation term using big $M$ constants. These terms deactivate the cut whenever a required charging station is not selected, thus ensuring that the value function cut only applies when the previous follower solution remains feasible. Let $M_{l}$ be a sufficiently large constant for each $l\in\mathcal{C}$, and define the set $\mathcal{C}^{\mathrm{f}}:=\{l\in\mathcal{C}:v_{l}^{\mathrm{f}}>0\}$, i.e., the stations that were used in the previous follower solution. The relaxed value function cuts become:
$L(\mathbf{y},\mathbf{v}(\mathbf{h}),\mathbf{x}(\mathbf{h}))\leq L(\mathbf{y},\mathbf{v}^{\mathrm{f}}(i),\mathbf{x}^{\mathrm{f}}(i))+\sum_{l\in\mathcal{C}}M_{l}(1-y_{l}),\quad\forall\,i\in n^{\mathbf{v}^{\mathrm{f}}}.$ (8)
@@
$y_{l}\in\{0,1\},$ $\quad\forall l\in\mathcal{C}$ (9c)
@@
$v_{l}=\sum_{\pi\in\Pi(\mathbf{y})}\delta_{l}^{\pi}h_{\pi},$ $\quad\forall l\in\mathcal{C}$ (9g)
@@
$L(\mathbf{y},\mathbf{v}(\mathbf{h}),\mathbf{x}(\mathbf{h}))\leq L(\mathbf{y},\mathbf{v}^{\mathrm{f}}(i),\mathbf{x}^{\mathrm{f}}(i))+\sum_{l\in\mathcal{C}}M_{l}(1-y_{l})$ $\quad\forall\,i\in n^{\mathbf{v}^{\mathrm{f}}}.$ (9j)

### v2

@@ 3.1 LL Problem (User Charging Decision)
In real networks, not all EV travel demands between OD pairs can be satisfied due to limited battery range and sparse charging infrastructure. This is an issue for travel over long distances, since usually the trips within a city can be completed without needing to charge along the route. As a result, for trips with long distances, certain trips become infeasible if no reachable path exists that either stays within the battery limit or includes a charging opportunity. This constraint motivates defining the set of feasible paths based on battery limitations, which acknowledges that some OD pairs may have zero flow in the UE solution. The feasible paths available to each user depend on the installed charging infrastructure $\mathbf{y}$.
@@
The LL problem thus jointly determines equilibrium path flows, link flows, and charging station demand, conditional on the UL decision $\mathbf{y}$. We consider a network $\mathcal{G}=(\mathcal{N},\mathcal{A})$: graph $\mathcal{G}$ is the tuple $(\mathcal{N},\mathcal{A})$ comprised of the set of nodes $\mathcal{N}$ and the set of links $\mathcal{A}$. Each link can be denoted by a pair of nodes $(i,j)$. Let $\mathcal{Z}\subseteq\mathcal{N}$ be the set of zones, which represent nodes that are the origins or destinations of the travelers. $\mathcal{C}\subseteq\mathcal{N}$ is the set of candidate locations for building charging infrastructure. A path is feasible for an EV if either the usage of the battery is less than the battery limit, assuming that users will start their trip with a full battery, or there is a charging station on the route where the user recharges the battery to reach the destination before the battery is depleted. Therefore, not all demand from an origin to a destination can be met, and it depends on the $\mathbf{y}$, which specifies where charging stations are located and consequently which paths are feasible. Let $\Pi_{rs}(\mathbf{y})$ be the set of feasible paths based on the charging stations location $\mathbf{y}$ between the origin $r\in\mathcal{Z}$ and destination $s\in\mathcal{Z}$. Let $\Pi(\mathbf{y)}=\cup_{(r,s)\in\mathcal{Z}^{2}}\Pi_{rs}(\mathbf{y})$ be the set of all feasible paths. Travel time is considered the travel cost. The travel time for the link $(i,j)$ is based on a convex function of traffic flow on this link, which is $t_{ij}(x_{ij})$, where $x_{ij}$ is the link $(i,j)$ flow. Similarly, the waiting time due to the congestion and delay at station $l$ is modeled as a convex function $t_{l}(v_{l})$ of the charging flow $v_{l}$. Before presenting the formulation, we summarize the key notation used throughout the model in Table 1.
Table 1: Summary of Notation
Notation Description
Sets
$\mathcal{G}=(\mathcal{N},\mathcal{A})$ Transportation network with nodes $\mathcal{N}$ and links $\mathcal{A}$
$\mathcal{Z}\subseteq\mathcal{N}$ Set of zones (origins and destinations)
$\mathcal{C}\subseteq\mathcal{N}$ Set of candidate locations for charging stations
$\Pi_{rs}(\mathbf{y})$ Set of feasible paths from origin $r$ to destination $s$ under design $\mathbf{y}$
$\Pi(\mathbf{y})$ Set of all feasible paths under design $\mathbf{y}$
Parameters
$D_{rs}$ Total potential demand between origin $r$ and destination $s$
$d_{rs}(\mathbf{y})$ Effective demand between $r$ and $s$ (depends on path feasibility)
$\delta_{ij}^{\pi}$ Binary indicator: 1 if path $\pi$ uses link $(i,j)$, 0 otherwise
$\phi_{l}^{\pi}$ Binary indicator: 1 if path $\pi$ includes charging at station $l$, 0 otherwise
$\gamma_{l}^{\pi}$ Charging time at station $l$ for a vehicle on path $\pi$
$c$ Charging price per unit of time
$\tau$ Value of time (converts charging duration to monetary cost)
$c^{\pi}$ Total charging cost for path $\pi$: $c^{\pi}=\sum_{l\in\mathcal{C}}(c+\tau)\cdot\gamma_{l}^{\pi}$
$t_{ij}(x_{ij})$ Travel time function on link $(i,j)$ as a function of flow
$t_{l}(v_{l})$ Waiting time function at station $l$ as a function of charging flow
$t_{ij}^{0}$ Free-flow travel time on link $(i,j)$
$t_{l}^{0}$ Base charging time at station $l$ (free-flow service time)
$C_{ij}$ Capacity of link $(i,j)$
$K_{l}$ Service capacity of station $l$
$C_{l}$ Installation cost at candidate location $l$
$B$ Total available budget for charging station installation
$p$ Charging revenue per unit of time
$w$ Weight for unmet demand penalty in objective function
$M,M_{l}$ Big-M constants
$\alpha,\beta$ Shape parameters for BPR function
$\kappa$ Capacity-cost proportionality constant: $K_{l}=\kappa\cdot C_{l}$
$B^{\max}$ Full battery capacity
$b_{l}^{\pi}$ Arrival battery level at station $l$ for a vehicle on path $\pi$
$\rho$ Charging rate (time required to charge one unit of battery)
Decision Variables
$y_{l}$ Binary variable: 1 if charging station is built at location $l$, 0 otherwise
$h_{\pi}$ Flow on path $\pi$
$x_{ij}$ Flow on link $(i,j)$
$v_{l}$ Total charging flow at station $l$
The full formulation of this convex UE problem is given below.
$\min_{\mathbf{x},\mathbf{h},\mathbf{v}}\quad$ $L(\mathbf{y,x(h),v(h)})=\sum_{(i,j)\in\mathcal{A}}\int_{0}^{x_{ij}}t_{ij}(x)\,dx+\sum_{l\in\mathcal{C}}\int_{0}^{v_{l}}t_{l}(v)\,dv+\sum_{\pi\in\Pi(\mathbf{y})}c^{\pi}h_{\pi}$ (1a)
@@
$v_{l}=\sum_{\pi\in\Pi(\mathbf{y})}\phi_{l}^{\pi}h_{\pi},$ $\quad\forall l\in\mathcal{C}$ (1e)
@@
The LL problem (1) constitutes a modified traffic assignment problem (MTAP) that accounts for the unique cost structure faced by EV users. In contrast to classical TAP formulations, which only consider travel time as the cost, our model includes travel time, station congestion, and path-dependent charging costs, which indicates EV users must also select charging stations along their routes. This extension is reflected in the objective function, which includes three terms: link travel time, station congestion cost, and charging cost based on each path’s energy requirements. Furthermore, two additional constraints distinguish this model from standard TAP: (i) a station activation constraint (1b) ensures no users can charge at the location $l$ unless a charging facility is built there. In this constraint, $y_{l}$ comes from UL and is fixed here, and (ii) a flow conservation constraint that determines the amount of charging demand at each station based on users’ path choices. These modifications are grounded in the assumption that EV travelers behave rationally under perfect information and choose paths that minimize their total travel and charging cost.
We will name this problem MTAP ($\mathbf{y}$), and let $L(\mathbf{y},\mathbf{x}(\mathbf{h}),\mathbf{v}(\mathbf{h}))$ be the objective function value. To achieve the travel time between $r$ and $s$, we need to sum up all the links’ travel time between $r$ and $s$, which are in the user’s path. The link flows, and the path flows, denoted as $h_{\pi}$, are closely related. Let $\delta^{\pi}_{ij}$ denote whether link $(i,j)$ is used by path $\pi$, so $\delta^{\pi}_{ij}=0$ if path $\pi$ does not use link $(i,j)$, and $\delta^{\pi}_{ij}=1$ if it does. Based on constraint (1c), link flows are the sum of all the path flows that use that link.
@@
Let $v_{l}$ be the total flow of vehicles charging at station $l$ per unit time. We can only have a positive value for $v_{l}$ if the decision is to build a charging station there. Constraint (1b) shows this linking relation between $v_{l}$ and $y_{l}$. So if $y_{l}$ is 0, then $v_{l}$ must be 0 too. In constraint (1e), $\phi^{\pi}_{l}\in\{0,1\}$ indicates whether a vehicle traveling along path $\pi$ charges at station $l$, and $h_{\pi}$ is the flow on that path. Therefore, $v_{l}$ is equal to the total flow of vehicles that stop to charge at station $l$. To capture congestion effects at charging stations, we model the waiting and service time as a flow-dependent function $t_{l}(v_{l})$, where higher demand leads to increased delays. This reflects the realistic scenario where stations with limited capacity experience congestion as more users arrive.
The third term in the objective function captures the path-dependent charging cost. Let $\gamma^{\pi}_{l}$ denote the amount of time (in hours) that a vehicle traveling along path $\pi$ must spend charging at station $l$, which depends on the battery level upon arrival. The total charging cost for path $\pi$ is defined as $c^{\pi}=\sum_{l\in\mathcal{C}}(c+\tau)\cdot\gamma_{l}^{\pi}$, where $c$ is the charging price per unit of time and $\tau$ is value of time which converts waiting time to recharge to equivalent cost. This formulation ensures that vehicles arriving with different battery states pay different amounts based on how much they actually need to charge.
The constraints (1g) and (1f) require the non-negativity for the flow of the charging station and the flow of the path. The first term in the objective function of problem (1) corresponds to the classical Beckmann transformation, which is a potential function whose minimizer satisfies the UE conditions with respect to travel time. The second term incorporates the station congestion cost through the integral $\int_{0}^{v_{l}}t_{l}(v)dv$, where $t_{l}(v)$ is a convex, monotonically increasing function representing the marginal waiting time at station $l$ under flow $v$. This formulation ensures that as station utilization increases, users experience longer waiting times, which is reflected in their path choices at equilibrium. The third term $\sum_{\pi}c^{\pi}h_{\pi}$ is linear in path flows and captures the actual energy cost paid by users and the time value they spend charging, which varies by path based on battery consumption before reaching the charging station.
@@ 3.2 Bi-level Problem
It is important to note that the revenue term and the unmet demand penalty term capture fundamentally different aspects of the planning problem and do not represent double counting. The revenue term $\sum_{l\in\mathcal{C}}\int_{0}^{v_{l}}p\,dv$ represents the actual charging revenue collected from EV users who successfully complete their trips and charge at the installed stations. In contrast, the unmet demand penalty captures the social cost of travelers who cannot complete their trips at all due to insufficient charging infrastructure. These two terms are mutually exclusive for any given demand: users either contribute to revenue (if they can travel and a feasible path exists) or to the unmet demand penalty (if no feasible path exists), but never both.
Furthermore, the specific objective function presented here represents one particular planning scenario. Our algorithmic framework is general and can accommodate alternative upper-level objectives depending on stakeholder interests. For instance, a private operator focused solely on profit could set $w=0$, while a public agency prioritizing accessibility could increase $w$ or adopt entirely different objective functions such as minimizing total system travel time or maximizing network coverage.
@@ 4.1 High Point Relaxation (Global Lower Bound)
We will derive the HPR problem by considering only the follower’s constraints without taking into account its objective function and relaxing the binary variable $y$ to be continious.
@@
$0\leq y_{l}\leq 1,$ $\quad\forall l\in\mathcal{C}$ (4c)
@@
$v_{l}=\sum_{\pi\in\Pi(\mathbf{y})}\phi_{l}^{\pi}h_{\pi},$ $\quad\forall l\in\mathcal{C}$ (4g)
@@ 4.2 Value Function Cuts
$0\leq y_{l}\leq 1,$ $\quad\forall l\in\mathcal{C}$ (6c)
@@
$v_{l}=\sum_{\pi\in\Pi(\mathbf{y})}\phi_{l}^{\pi}h_{\pi},$ $\quad\forall l\in\mathcal{C}$ (6g)
@@
$L(\mathbf{y},\mathbf{v}(\mathbf{h}),\mathbf{x}(\mathbf{h}))\leq L(\mathbf{y},\mathbf{v}^{\prime}(\mathbf{h}^{\prime}),\mathbf{x}^{\prime}(\mathbf{h}^{\prime})),$ $\quad\forall\,\mathbf{h}^{\prime}\in\mathcal{H}(\mathbf{y})$ (6j)
Constraint (6j) represents the value function cuts. $\mathcal{H}(\mathbf{y})$ represents the set of feasible path flow vectors $\mathbf{h}^{\prime}$ under the network design $\mathbf{y}$, and $\mathbf{v}^{\prime}(\mathbf{h}^{\prime})$, $\mathbf{x}^{\prime}(\mathbf{h}^{\prime})$ are the corresponding charging and link flows. In particular, for each $\mathbf{h}^{\prime}\in\mathcal{H}(\mathbf{y})$, the derived aggregate flows $\mathbf{v}^{\prime}(\mathbf{h}^{\prime})$ and $\mathbf{x}^{\prime}(\mathbf{h}^{\prime})$ respect $v_{l}^{\prime}=0$ if $y_{l}=0$ for all $l\in\mathcal{C}$, so that no site without an established charging station is assigned any charging demand. By restricting to flow configurations consistent with the given $\mathbf{y}$, this formulation avoids infeasible cases such as $v_{l}^{\prime}>0$ when $y_{l}=0$.
@@
It is sufficient to show equality of the feasible regions, since the objective functions in both problems are identical. Suppose $(\mathbf{y},\mathbf{v},\mathbf{x})$ is feasible for problem (6). This means $(\mathbf{y},\mathbf{v},\mathbf{x})$ satisfies the value function cuts in constraint (6j). Therefore, under the same $\mathbf{y}$, the $L(\mathbf{y},\mathbf{v}(\mathbf{h}),\mathbf{x}(\mathbf{h}))$ is less than or equal to the $L(\mathbf{y},\mathbf{v}^{\prime}(\mathbf{h}^{\prime}),\mathbf{x}^{\prime}(\mathbf{h}^{\prime}))$ for all $(\mathbf{v}^{\prime}(\mathbf{h}^{\prime}),\mathbf{x}^{\prime}(\mathbf{h}^{\prime}))$ that are dervived form $\mathbf{h}^{\prime}\in\mathcal{H}(\mathbf{y})$. This directly means $(\mathbf{v},\mathbf{x})$ minimizes $L(\mathbf{y},\mathbf{v}(\mathbf{h}),\mathbf{x}(\mathbf{h}))$ given the decision $\mathbf{y}$. Thus, we can ensure the solution is feasible for the original bi-level problem (6). Conversely, consider any point $(\mathbf{y},\mathbf{v},\mathbf{x})$ feasible for the original bi-level problem (6). By the definition of LL optimality, $(\mathbf{v},\mathbf{x})$ minimizes $L(\mathbf{y},\mathbf{v}(\mathbf{h}),\mathbf{x}(\mathbf{h}))$ given the decision $\mathbf{y}$. Therefore, constraint (6j) is trivially satisfied, implying feasibility for problem (6). Hence, both problems have identical feasible sets and are thus equivalent. ∎
@@
[…]
$L(\mathbf{y},\mathbf{v}(\mathbf{h}),\mathbf{x}(\mathbf{h}))\leq L(\mathbf{y},\mathbf{v}^{\mathrm{f}}(i),\mathbf{x}^{\mathrm{f}}(i))+\sum_{l\in\mathcal{C}}M_{l}(1-y_{l}),\quad\forall\,i\in\mathcal{F}$ (8)
@@
$0\leq y_{l}\leq 1,$ $\quad\forall l\in\mathcal{C}$ (9c)
@@
$v_{l}=\sum_{\pi\in\Pi(\mathbf{y})}\phi_{l}^{\pi}h_{\pi},$ $\quad\forall l\in\mathcal{C}$ (9g)
@@
$L(\mathbf{y},\mathbf{v}(\mathbf{h}),\mathbf{x}(\mathbf{h}))\leq L(\mathbf{y},\mathbf{v}^{\mathrm{f}}(i),\mathbf{x}^{\mathrm{f}}(i))+\sum_{l\in\mathcal{C}}M_{l}(1-y_{l})$ $\quad\forall\,i\in\mathcal{F}.$ (9j)